OpenStudy (anonymous):

what multiplies to equal 168 but adds up to -34

OpenStudy (campbell_st):

so let the numbers be x and y looking at the question they would both have to be negative numbers so you know \[x \times y = 168\] and \[x + y = -34\] so use some substitution take the 2nd equation and make x the subject \[x = -34 - y\] and substitute into the 1st equation \[(-34 -y) \times y = 168\] which becomes \[-34y - y^2 = 168....or..... y^2 + 34y + 168 = 0\] which factors to (y + 28)(y + 6) = 0 so the solutions are -28 and - 6 hope it helps

OpenStudy (anonymous):

wow, thts a lot more than i expected, thanks